# Think-at-Hard: Selective Latent Iterations to Improve Reasoning Language Models

### v1

@@ A.2.4 Iteration Decider Analysis
Sensitivity Analysis. We conduct additional experiments to analyze the sensitivity of iteration decisions by perturbing the oracle iteration decider with different types of noise. We investigate the differing impacts of overthink and underthink, where overthink refers to cases where the decider incorrectly signals “continue,” and underthink corresponds to incorrect “stop” signals. The results in Table 10 demonstrate that the performance of latent reasoning models is highly sensitive to the accuracy of iteration decisions. We quantify these impacts by fitting a linear model to the data:
$\text{acc.}=-1.41\times\#\text{underthink}-2.73\times\#\text{overthink}+0.81$
Further analysis reveals that the primary factor contributing to the performance gap between TaH and its oracle variant, TaH-oracle, is inaccurate iteration decisions, with overthink being the dominant source of errors.
Table 10: Oracle model performance with overthinks and underthinks. All values are reported in percentages.
False rate Underthink Overthink Accuracy
0 0 0 80.0
@@
Table 11: Parameter breakdown of Standard, TaH and TaH+. Counts are reported using M (million) and B (billion).
Param. Method Backbone LoRA Iter. Decider Total
0.6B Standard 596M 596M
TaH 580M 10M 5M 595M
TaH+ 596M 10M 5M 611M
1.7B Standard 1.72B 1.72B
TaH 1.67B 34M 18M 1.72B
TaH+ 1.72B 34M 18M 1.77B
A.2.5 Attention Pattern Analysis
@@
Figure 8(a) further presents the total attention scores assigned to keys in the first iteration. It can be seen that the first layer tends to focus more on keys from the second iteration. Different layers also exhibit varying attention behaviors.
(a) TaH mean and standard deviation of attention weights (key from iteration 1) across layers in iteration 2.
(b) Validation perplexity for different training schemes.
A.2.6 Design Choice Exploration
Figure 8(b) shows validation perplexity (ppl) across training schemes. TaH achieves lower perplexity than iter. decider and token+latent. Although dynamic attains the lowest perplexity, it collapses on downstream tasks, often generating infinite-loop outputs.
A.3 Implementation Details
A.3.1 Duo-causal attention implementation
@@
This realizes the accessible set in Equation 5, allowing deeper iterations to access all shallower iterations while preserving positional causality.
(2) Two-dimensional causal mask. For a query $(i,d)$, a key $(j,k)$ is attendable iff $j\leq i$ and $k\leq d$. We implement this as an additive attention mask with 0 for allowed entries and -$\infty$ otherwise, enforcing positional and iteration causality jointly. When $d=1$ for all tokens, the rule reduces to standard causal attention.
@@
A.4 Addtional Related Work
@@
Layer Skipping. Layer skipping aims to accelerate LLM inference by dynamically bypassing certain layers for specific tokens. Some methods use a learnable module to make real-time skipping decisions. MoD (raposo2024mixture) uses a top-k router to select a subset of tokens for processing, while FlexiDepth (luo2025adaptive) uses a plug-in router to determine whether a layer should be bypassed. Others use a fixed strategy to skip layers. SkipDecode (del2023skipdecode) enforces a monotonically decreasing number of active layers during generation. However, these methods still require loading the entire model’s parameters, resulting in a large memory access overhead. Instead of skipping some layers, TaH adds computational depth by allowing core tokens to undergo multiple refinement iterations. This approach provides greater computational depth without increasing the model’s parameter count.
A.5 Limitations and Future Work
Comparison with Official Qwen3 Models. Official Qwen3 models are trained on different data distributions and scales, and use different training procedures, including on-policy distillation (yang2025qwen3). By contrast, our models use SFT only on limited, publicly accessible data. Consequently, performance may differ between the two.
@@ A.5 Limitations and Future Work
A.6 Use of Large Language Models (LLMs)

### v2

@@ A.3.5 Iteration Decision Error
We analyze how iteration decision accuracy affects TaH’s end-to-end response quality, since iteration decider will not be perfect as shown in Figure 9. To this end, we randomly inject errors into the oracle iteration-decider predictions at different rates. Formally, we denote the original oracle prediction as the label $l\in\{0,1\}$ and the altered prediction as the output $o\in\{0,1\}$. We define the iter. error as the total proportion of deliberately introduced errors:
$\text{iter. error}=P(l\neq o)=\underbrace{P(l=1,o=0)}_{\textit{underthink rate}}+\underbrace{P(l=0,o=1)}_{\textit{overthink rate}}.$ (10)
We further distinguish the impacts of overthinking and underthinking. Here, overthinking refers to cases where the decider incorrectly signals continue, while underthinking corresponds to cases where it incorrectly signals stop. Table 18 shows how TaH’s MATH100 accuracy varies with different iteration error rates. We quantify these effects by fitting a linear model to the data:
$\text{accuracy}=-1.41\times\text{underthink rate}-2.73\times\text{overthink rate}+0.81.$
This analysis indicates that inaccurate iteration decisions are the main factor behind the performance gap between TaH and its oracle variant, with overthinking being the dominant source of performance gaps.
Table 18: TaH performance under different iteration-decider error rates. All values are reported in percentages.
Iter. Error (%) Underthink (%) Overthink (%) MATH100 Accuracy (%)
0.0 0.0 0.0 80.0
2.8 2.8 0.0 78.0
@@
22.1 0.0 22.1 21.6
A.3.6 Duo-causal Attention Pattern
@@
Figure 12 further presents the total attention scores assigned to keys in the first iteration. It can be seen that the first layer tends to focus more on keys from the second iteration. Different layers also exhibit varying attention behaviors.
Figure 11: TaH mean and standard deviation of attention weights (key from iteration 1) across layers in iteration 2.
Figure 12: Validation perplexity for different training schemes.
A.4 Implementation Details
A.4.1 Duo-causal attention implementation
[…]
Figure 13 illustrates the implementation of duo-causal attention, with the formal definitions provided below.
@@ A.4.1 Duo-causal attention implementation
This realizes the accessible set in Equation 6, allowing deeper iterations to access all shallower iterations while preserving positional causality. The KV cache is managed by iteration depth during decoding, as shown in Figure 13(b). The fragmented KV-cache management strategy is standard in existing LLM serving systems (Kwon et al., 2023; Zheng et al., 2024).
(2) Two-dimensional causal mask. For a query $(i,d)$, a key $(j,k)$ is attendable iff $j\leq i$ and $k\leq d$. We implement this as an additive attention mask with 0 for allowed entries and -$\infty$ otherwise, enforcing positional and iteration causality jointly. Figure 13(c) visualize the landscape of the duo-causal attention mask. When $d=1$ for all tokens, the rule reduces to standard causal attention.
@@
A.5 Additional Related Work
Signal-guided Control. These methods keep reasoning in token space but steer computation by inserting control tokens. They add filler tokens (e.g., dots) (Pfau et al., 2024), learnable [PAUSE] tokens (Goyal et al., 2024; Kim et al., 2025) for extra compute during decoding. They are lightweight but constrained to discrete-token interventions with limited latent control.
Latent Optimization. These methods perform autoregressive reasoning directly in internal representations, emitting little or no intermediate text. They distill CoT into continuous embeddings via progressive replacement (Hao et al., 2024; Cheng & Van Durme, 2024), hidden-state alignment (Su et al., 2025; Liu et al., 2024), or logit-weighted embeddings (Zhang et al., 2025). While efficient, these methods sacrifice interpretability, with training-based ones requiring heavy mitigation from verbal LLMs.
@@ A.5 Additional Related Work
Layer Skipping. Layer skipping aims to accelerate LLM inference by dynamically bypassing certain layers for specific tokens. Some methods use a learnable module to make real-time skipping decisions. MoD (Raposo et al., 2024) uses a top-k router to select a subset of tokens for processing, while FlexiDepth (Luo et al., 2025) uses a plug-in router to determine whether a layer should be bypassed. Others use a fixed strategy to skip layers. SkipDecode (Del Corro et al., 2023) enforces a monotonically decreasing number of active layers during generation. However, these methods still require loading the entire model’s parameters, resulting in a large memory access overhead. Instead of skipping some layers, TaH adds computational depth by allowing core tokens to undergo multiple refinement iterations. This approach provides greater computational depth without increasing the model’s parameter count.
A.6 Limitations and Future Work
Comparison with Official Qwen3 Models. Official Qwen3 models are trained on different data distributions and scales, and use different training procedures, including on-policy distillation (Yang et al., 2025). By contrast, our models use SFT only on limited, publicly accessible data. Consequently, performance may differ between the two.
@@ A.6 Limitations and Future Work
A.7 Use of Large Language Models (LLMs)